OpenStudy (anonymous):

can you check my work? In () are the ones i chose 1.Given the function f(x) = 2(x + 8), find x if f(x) = 12 (2) 40 −2 14 #2.Given the function g(x) = −4x + 5, find g(−4). −11 21 (2.25) 11 #3. If h(x) = −2x − 10, find h(−4). −2 −18 (−3) −16

OpenStudy (anonymous):

[3.06] If f(x) = 2x2 − 10, find f(5). 90 10 7.5 40 I also dont know the answer to this one

OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 22. −36 (−7) 52 −15

OpenStudy (anonymous):

number 2 is wrong I think : \[g(-4) = -4 \times -4 +5 = 16 + 5 = 21\]

OpenStudy (anonymous):

on number 3 you got the same mistake as number 2 : \[h(-4) = -4 \times -2 - 10 = 8 - 10 = -2\]

OpenStudy (anonymous):

see, first you solve the multplication, remember that a negative number times a negative number is a positive number then you subtract.

OpenStudy (anonymous):

okay so g(x) = −4x + 5, find g(−4) do I plug in -4 in both the x .

OpenStudy (anonymous):

you plug -4 in where you got a x in the function. g(-4) means you are going to find what number the function g will give you if you plug -4 in place of x

OpenStudy (anonymous):

the last one ou posted is correct and the one before that I didnt understand the function

OpenStudy (anonymous):

its fine i got it thank you!